the radii of the two columns in U tube are r1 and r2
when a liquid of density p (rho) angle of contact = 0 degree is filled in it the level difference of liquid in two arms is h. the surface tension of liquid is

Grade:12th pass

1 Answers

Eshan
5 years ago
Drop in pressure due to formation of meniscus in first column=\dfrac{2T}{r_1}
Drop in pressure due to formation of meniscus in first column=\dfrac{2T}{r_2}

Hence the heights of liquids in the two columns differ to balance the excess drop in pressure in one of hte arms.

Hence\rho gh=2T(\dfrac{1}{r_1}-\dfrac{1}{r_2})
\implies T=\dfrac{\rho ghr_1r_2}{2(r_2-r_1)}